OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system. 6x – 5y = 4 9x + 4y = –17 Thanks !

OpenStudy (anonymous):

kelliegirl13, you got this? :)

OpenStudy (anonymous):

The elimination method requires you to " eliminate " either the x's or the y's to find the other letter. To do this, you sometimes have to multiply one or both of the equations by a number so that the x's or the y's cancel each other out. We will begin now... 6x - 5y = 4 --->(4)6x - 5y = 4 9x + 4y = - 17 -->(5) 9x + 4y = - 17 -------------- 24x - 20y = 16 (result of multiplying by 4) 45x + 20y = - 85 (result of multiplying by 5) ---------------add 69x + 0 = - 69 69x = - 69 x = -69/69 = - 1 we have found x, so now we must substitute -1 in for x in either of the original equations. 6x - 5y = 4 6(-1) - 5y = 4 -6 - 5y = 4 -5y = 4 + 6 -5y = 10 y = 10/-5 = -2 lets check our answers now... 9x + 4y = - 17 9(-1) + 4(-2) = -17 -9 - 8 = - 17 -17 = -17 (correct) ANSWER: x = -1 and y = -2 or (-1,-2) Any questions ?

OpenStudy (anonymous):

Nope, i'm good! Thanks!

OpenStudy (anonymous):

good to hear...glad I could help :)